OpenStudy (anonymous):

Simplify:24 divided by 1/2+5/8*6/7

OpenStudy (anonymous):

5/8 * 6/7 = 30/ 56 = 15 / 28 + 14/ 28 = 29/30 24 x 30 / 29 = answer. I dont have a calculator right now.

OpenStudy (campbell_st):

when dividing by a fraction, invert and multiply the first part of the question is asking how many halves in 24 when multiplying fractions... multiply the numerators and the denominators \[x \div \frac{a}{b} = x \times \frac{b}{a}\] and \[\frac{a}{b} \times \frac{c}{d} = \frac {a \times c}{b \times d}\]